# Inference for concave distribution functions under measurement error

### v1

@@ 5.2 About least favorable hypothesis
A common way to calibrate tests without bootstrap is to exhibit a least favorable null distribution $F_{0}$, i.e., a function satisfying the null hypothesis such that, under the null, $T_{n}(F_{n}^{d})$ is asymptotically stochastically smaller than $T_{n0}$ where $T_{n0}$ denotes the statistic obtained when the true underlying function equals $F_{0}$. Since $F_{0}$ is known, one can easily simulate $T_{n0}$ and compute its $1-\gamma$ quantile using Monte Carlo. Then, the test that rejects the null hypothesis if $T_{n}(F_{n}^{d})$ exceeds this quantile has (asymptotic) level $\gamma$, but it has smaller power than that of the test calibrated using bootstrap. Anyway, this kind of calibration is appealing since it is easier to implement than the bootstrap, so it would be interesting to see whether or not it could be implemented in our case. Unfortunately, we are not able to exibit a least favorable hypothesis in our context, and we explain why below.
Let us start with a few literature. Testing concavity (or convexity) of a cumulated hazard rate on a given bounded interval under possible censoring is considered in [8]. Testing monotonicity of a regression function, which is equivalent to testing concavity or convexity of its primitive, on a given bounded interval is considered in [10]. In both cases, the test statistic is built using least concave majorants, and affine functions are least favorable for concavity (or convexity). This is proved using the facts that the distribution of the considered estimator in [8] after centering does not depend on the function $F$ on which the test is performed, and that in [10] has the same distribution as $G_{n}\circ F$ where $G_{n}$ is a process whose distribution does not depend on $F$. Hence in both cases the way the distribution of the estimator depends on $F$ is quite simple.
Based on the literature, it is natural to wonder if affine functions are least favorable in our context. Unfortunately, it seems that it is not the case.
@@
Indeed, all uniform distributions are affine on a bounded support, which makes them all good candidates for being least favourable. However, uniform distributions just differ in terms of scale parameters. In the classic framework, where there is no measurement error, scaling does not make any difference, so one can conveniently choose the uniform on $[0,1]$. However, in our setting this is not possible. If we rescale $X$, this corresponds to rescaling $\epsilon$, that is, changing the noise-to-signal ratio, which has a crucial impact on the estimation accuracy, and consequently, on the power of the test. Since the noise-to-signal ratio is generally unknown, it is not possible to know which uniform distribution is the least favourable in a given case. Moreover, even assuming that the variance of $X$ is known, we are not able to prove that affine functions are least favorable because the way the distribution of our estimator depends on our function $F$ is very intricate. Indeed, the limit covariance $\Sigma_{s,t}$ depends on $f_{Y}=f_{\varepsilon}\star f$ which cannot be expressed in terms of $f_{\varepsilon}\star f_{0}$, with $f_{0}$ indicating the density of the uniform distribution on the support of $f$.
A way to circumvent the absence of a least favorable hypothesis, at the cost of a slight loss of power, is to reject the null hypothesis if $T_{n}(F_{n}^{d})$ exceeds $v_{n}$, an arbitrary sequence of positive numbers that tends to infinity as $n\to\infty$ with $\sqrt{n}v_{n}$ that tends to zero. It follows from theorem 5 that $T_{n}(F_{n}^{d})$ is bounded in probability under the null hypothesis while it diverges to infinity at the rate $\sqrt{n}$ under the alternative, whence the above test has asymptotic level equal to zero and asymptotic power equal to one at any point in the alternative. The smaller $v_{n}$ is, the larger the power. One can choose for instance $v_{n}=\log n$.
@@ 6.1 Performance of $\mathcal{M}\hat{F}_{n}^{d}$ in $\mathcal{H}_{0}$
This section evaluates the performance of the constrained estimator $\mathcal{M}\hat{F}_{n}^{d}$ in terms of its mean squared error (MSE). The objective is to determine whether the constrained estimator $\mathcal{M}\hat{F}_{n}^{d}$ provides improvements over its unconstrained counterpart $\hat{F}_{n}^{d}$, when $F$ is concave. To this end, we analyze the ratio of the MSE of $\mathcal{M}\hat{F}_{n}^{d}$ and $\hat{F}_{n}^{d}$ computed at different quantiles, based on $500$ simulation runs. Here, $\mathcal{M}\hat{F}_{n}^{d}$ outperforms $\hat{F}_{n}^{d}$ iif the MSE ratio is smaller than one.
@@
For simplicity, we restrict our attention to the case where the measurement error follows a Laplace distribution, with mean 0 and standard deviation $\sigma_{\varepsilon}$. The standard deviation is chosen to achieve a specific noise-to-signal ratio (NSR), following [6, p. 483], defined as
@@ 6.3 The performance of the test when the error alters the shape
In some situations, the presence of measurement error affects the shape of the convolution substantially. For example, when $X$ follows the mixture distribution $X\sim 0.2\,W(3,1)+0.8\,B(0.5,0.75),$ then $F$ satisfies $\mathcal{H}_{1}$. However, after convolution with measurement error, the distribution function of $X+\varepsilon$ may fall into $\mathcal{H}_{0}$, especially when the NSR is large. In such cases, it is essential to apply the proposed test in order ton discover the shape of the true underlying distribution $F$.
@@
Appendix A Proofs for Sections 2 and 3
A.1 Proof of Lemma 1
Arguing as in the proof of Theorem 3.1 in [19] with $M=1/h$ proves that
and that
$\mathbb{E}\|f_{n}^{d}-f\|_{\infty}=o(1).$
It then follows from Markov’s inequality that
$\|f_{n}^{d}-f\|_{\infty}=o_{p}(1).$
Moreover, for all $t\geq 0$, one has
where $\lim_{t\to\infty}F(t)=1$. Therefore,
$\lim_{t\to\infty}F_{n}^{d}(t)=1+o_{p}(1),$ (17)
A.2 Proof of Theorem 1
$\mathcal{M}F_{n}(y_{n}+t)\leq\mathcal{M}F_{n}(y_{n})-u_{n}t$
We turn to the proof of the second claim. For all $t\geq 0$, one has
where by assumption, the supremum on the right-hand side tends to zero. Then, for every $t\in\mathbb{R}^{+}$, $F_{n}(t)$ converges to $F(t)$. Furthermore, the convergence is uniform over $\mathbb{R}^{+}$ by applying similar arguments as in the proof of Lemma 2.11 in [32], using that $F$ is monotone (by non-negativity of $f$), continuous and bounded. Then, it follows from Marshall’s inequality [14, Exercise 3.1] that
$\|\mathcal{M}F_{n}-\mathcal{M}F\|_{\infty}\leq\|F_{n}-F\|_{\infty},$
whence the left-hand side converges to zero.
and for all $u<v$. Morever, the triangle inequality yields that
$\leq$ $\|f_{n}-f\|_{\infty}=o(1).$
Hence,
$=$ $\inf_{v>x}\sup_{u<x}\left(\frac{1}{v-u}\int_{u}^{v}f(t)dt\right)+o(1)$
$=$ $f_{0}(x)+o(1).$
For $x=0$ the min max formulas become
Hence,
$f_{n0}(0)$ $=$ $\inf_{v>0}\left(\frac{1}{v}\int_{0}^{v}f_{n}(t)dt\right)$
$=$ $f_{0}(0)+o(1).$
A.3 Proof of Corollary 1
A.4 Proof of Theorem 4
Combining with the functional delta method and Theorem 3 yields
Appendix B Proof of Theorem 5
$T_{n}(F_{n}^{d})=\sqrt{n}\|\mathcal{W}F_{n}^{d}-\mathcal{W}F\|_{\infty}.$ (18)
Now, assume that $F$ is not concave on $\mathbb{R}^{+}$ and define
$A:=\|F-\mathcal{M}F\|_{\infty},$
Using Marshall’s inequality [14, Exercise 3.1] on gets
$\geq$ $A-2\|F-F_{n}^{d}\|_{\infty}.$
Therefore,
Appendix C Generalization of the Donsker theorem in [30] to a bootstrap setting
@@
with $f_{X}$ being the probability density function of the $X_{i}$s, where the special case $\zeta:=\mathds{1}_{(-\infty,0]}$ leads to the estimation of the distribution function of the $X_{i}$s. Theorem 1 in [30] states that $\sqrt{n}(\widehat{\vartheta_{t}}-\vartheta_{t})_{t\in\mathbb{R}}$ converges in law to a centered Gaussian process $\mathbb{G}$ in $\ell^{\infty}(\mathbb{R})$, whose covariance depends on the functional $\zeta$ and on the distribution of the errors $\varepsilon_{j}$. Here,

### v2

@@ 5.3 About least favorable hypothesis
A common way to calibrate tests without bootstrap is to exhibit a least favorable null distribution $F_{0}$, i.e., a function satisfying the null hypothesis such that, under the null, $T_{n}(F_{n}^{d})$ is asymptotically stochastically smaller than $T_{n0}$ where $T_{n0}$ denotes the statistic obtained when the true underlying function equals $F_{0}$. Since $F_{0}$ is known, one can easily simulate $T_{n0}$ and compute its $1-\gamma$ quantile using Monte Carlo. Then, the test that rejects the null hypothesis if $T_{n}(F_{n}^{d})$ exceeds this quantile has (asymptotic) level $\gamma$, but it has smaller power than that of the test calibrated using bootstrap. Anyway, this kind of calibration is appealing since it is easier to implement than the bootstrap, so it would be interesting to see whether or not it could be implemented in our case.
@@
When testing concavity of a distribution function using the $\ell_{\infty}$-distance between the empirical distribution function and its least concave majorant as a test statistic, based on an i.i.d. sample from that distribution, the uniform is known to be least favorable, see [25]. Hence, it is natural to wonder if uniform distributions are least favorable in our context. Unfortunately, it seems that it is not the case. Indeed, uniform distributions just differ in terms of scale parameters. In the classic framework, where there is no measurement error, scaling does not make any difference, so one can conveniently choose the uniform on $[0,1]$. However, in our setting this is not possible. If we rescale $X$, this corresponds to rescaling $\epsilon$, that is, changing the noise-to-signal ratio, which has a crucial impact on the estimation accuracy, and consequently, on the power of the test. Since the noise-to-signal ratio is generally unknown, it is not possible to know which uniform distribution is the least favourable in a given case. Moreover, even assuming that the variance of $X$ is known, we are not able to prove that affine functions are least favorable because the way the distribution of our estimator depends on our function $F$ is very intricate. Indeed, the limit covariance $\Sigma_{s,t}$ depends on $f_{Y}=f_{\varepsilon}\star f$ which cannot be expressed in terms of $f_{\varepsilon}\star f_{0}$, with $f_{0}$ indicating the density of the uniform distribution on the support of $f$.
A simple way to circumvent the absence of a least favorable case, at the cost of conservativeness, is to reject the null hypothesis whenever $T_{n}(F_{n}^{d})$ exceeds a deterministic threshold $v_{n}$, where $v_{n}\to\infty$ and $v_{n}=o(\sqrt{n})$ as $n\to\infty$. For instance, one may take $v_{n}=\log n$. Under the null hypothesis, Theorem 5 establishes the limit distribution of the test statistic, which is almost surely finite, and consequently $T_{n}(F_{n}^{d})=O_{P}(1)$. Then, $P(T_{n}(F_{n}^{d})>v_{n})\to 0$, hence, the resulting test has asymptotic level equal to zero and is conservative. Under any fixed alternative, the test statistic diverges to infinity at rate $\sqrt{n}$. Since $v_{n}=o(\sqrt{n})$, it follows that the test is consistent and has asymptotic power equal to one at every point in the alternative.
@@ 6.1 Performance of $\mathcal{M}\hat{F}_{n}^{d}$ in $\mathcal{H}_{0}$
This section evaluates the performance of the constrained estimator $\mathcal{M}\hat{F}_{n}^{d}$ in terms of its mean squared error (MSE). The objective is to determine whether the constrained estimator $\mathcal{M}\hat{F}_{n}^{d}$ provides improvements over its unconstrained counterpart $\hat{F}_{n}^{d}$, when $F$ is concave. To this end, we analyze the ratio of the MSE of $\mathcal{M}\hat{F}_{n}^{d}$ and $\hat{F}_{n}^{d}$ computed at different quantiles, based on $500$ simulation runs. Here, $\mathcal{M}\hat{F}_{n}^{d}$ outperforms $\hat{F}_{n}^{d}$ when the MSE ratio is smaller than one.
@@
For simplicity, we restrict our attention to the case where the measurement error follows a Laplace distribution, with mean 0 and standard deviation $\sigma_{\varepsilon}$. The standard deviation is chosen to achieve a specific noise-to-signal ratio (NSR), following [10, p. 483], defined as
@@ 6.3 The performance of the test when the error alters the shape
In some situations, the presence of measurement error affects the shape of the convolution substantially. For example, when $X$ follows the mixture distribution $X\sim 0.2\,W(3,1)+0.8\,B(0.5,0.75),$ then $F$ satisfies $\mathcal{H}_{1}$. However, after convolution with measurement error, the distribution function of $X+\varepsilon$ may fall into $\mathcal{H}_{0}$, especially when the NSR is large. In such cases, it is essential to apply the proposed test in order to discover the shape of the true underlying distribution $F$.
@@
Appendix A Generalization of the Donsker theorem in [41] to a bootstrap setting
@@
with $f_{X}$ being the probability density function of the $X_{i}$s, where the special case $\zeta:=\mathds{1}_{(-\infty,0]}$ leads to the estimation of the distribution function of the $X_{i}$s. Theorem 1 in [41] states that $\sqrt{n}(\widehat{\vartheta_{t}}-\vartheta_{t})_{t\in\mathbb{R}}$ converges in law to a centered Gaussian process $\mathbb{G}$ in $\ell_{\infty}(\mathbb{R})$, whose covariance depends on the functional $\zeta$ and on the distribution of the errors $\varepsilon_{j}$. Here,